# Detection of a new GeV source in the outer region of the Coma cluster: a signature of external accretion shock ?

### v1

@@ 2.2 Morphological and spectral analysis
We examine whether there is any known counterpart at the position of the source W. Since most of the high-latitude gamma-ray background sources are active galactic nuclei (AGNs) and they are sources of radio emission, we search for possible radio sources associated with the source W. We find no radio sources within 99% containment radius of the source W (see Appendix B for more details).
@@ 4 Conclusions and Discussions
The IC emission could extend down to X-ray band as long as the power-law distribution of CR electrons extends to a minimum Lorentz factor of $\gamma_{\rm min}\sim 10^{3}$. Assuming $\alpha=-1.81\pm 0.21$ for the photon index, we expected a differential flux of $F_{\rm X}\sim 10^{-15}-10^{-13}{~{}\rm erg~{}cm^{-2}~{}s^{-1}}$ at X-ray band for the source W. Deep observations with NuSTAR, eROSITA and Einstein Probe could test this prediction.

### v2

@@ 2.2 Morphological and spectral analysis
We examine whether there is any known counterpart at the position of the source W. Since most of the high-latitude gamma-ray background sources are active galactic nuclei (AGNs) and they are sources of radio emission, we search for possible radio sources associated with the source W. We find no bright radio sources within 99% containment radius of the source W (see Appendix B for more details).
@@ 4 Conclusions and Discussions
The IC emission could extend down to X-ray band as long as the power-law distribution of CR electrons extends to a minimum Lorentz factor of $\gamma_{\rm min}\sim 10^{3}$. Assuming $\alpha=-1.81\pm 0.21$ for the photon index, we expected a differential flux of $F_{\rm X}\sim 10^{-14}-10^{-13}{~{}\rm erg~{}cm^{-2}~{}s^{-1}}$ at X-ray band for the source W. Deep observations with NuSTAR, eROSITA and Einstein Probe could test this prediction.
@@ Appendix B Search for radio sources in the error box of the source W
In our search for possible X-ray counterparts in the eROSITA DR1 catalog  [^7] https://erosita.mpe.mpg.de/dr1/erodat/, we found a X-ray point source, 1eRASS J124657.2+272806, within the $1\sigma$ positional error of the source W, with a flux of $1.2\times 10^{-13}\,{\rm erg\,cm^{-2}\,s^{-1}}$. Additionally, by comparing this region with data from the VLASS survey  [^8] https://science.nrao.edu/vlass, we note the presence of a radio point source, VLASS1QLCIR J124657.34+272759.6, within the $2\sigma$ positional uncertainty region of the eROSITA X-ray source, with a flux density of $1.4\,{\rm mJy}$ in the 2–4 GHz band. According to the radio-GeV flux correlation for Fermi blazars (Ghirlanda et al., 2011), the expected GeV flux of this radio source is insufficient to account for the GeV flux of source W, although we cannot completely rule out the possibility of the association between two sources considering the uncertainty in the radio-GeV correlation.